The product of two consecutive odd integers is 2 less than six times their sum. Can you find the two integers
Answer by drk(1908)

x(x+2) = 6*(x+x+2) - 2
distribute to get
x^2 + 2x = 12x + 12 - 2
set = 0 to get
x^2 - 10x - 10 = 0
There are no solutions to this. The closest we get is 11, 13